OpenStudy (anonymous):

A bag has 4 green marbles, 3 red marbles, and 3 yellow marbles. What is the probability that you pick a green marble, do not replace it, and pick another green marble?

OpenStudy (anonymous):

In the beginning, the probability of a green marble is 4/10 After you've picked the first one and do not replace it, the probability of a green marble is 3/9. So your statement would be: Probability of green and green = Probability of 1st green * probability of 2nd green = (4/10) * (3/9) = 12/90 = 0.1333

OpenStudy (anonymous):

@mubzz is that the same as 2/15?

OpenStudy (anonymous):

yep